OpenStudy (anonymous):

A lamp shade has a radius of 9 inches. What is the circumference of the top of the shade? Use 3.14 to approximate pi. Round your answer to the nearest whole number.

OpenStudy (anonymous):

c = circumference pi = 3.14 r = radius = 9 inches d = diameter c=pi*d The radius is half of the diameter, d = 2r. Then d = 2*9=18. plug into equation. c= pi*d = 3.14*18 round accordingly and you may want to include that the measurement you end up with is in inches.